Reggie is 5 years older than Lani. Five years ago the sum of their ages was 25. How old are they now?
Answer by stanbon(75887)

Reggie is 5 years older than Lani. Five years ago the sum of their ages was 25. How old are they now?

NOW DATA:
Lani's age = x ; Reggie's age = x+5
FIVE YEARS AGO DATA:
Lani age was = x-5 ; Reggie age was : x+5-5 = x
EQUATION based on FIVE YEARS AGO DATA:
sum of their ages was 25
x-5+x = 25
2x=30
x=15
Lani's age NOW is 15 yrs
Reggie's age NOW is 20 yrs
Cheers,
Stan H.